How much of a 7% alcohol solution should be mixed with how much of a 15% alcohol solution so that you get 20 gallons of a 12% alcohol solution?
Answer by mathstutor494(120)

Let x be the number of gallons of 15% alcohol
Therefore
number of gallons of 7% alcohol = 20-x
x*0.15+(20-x)*0.07=20*0.12
0.15x+1.4-0.07x = 2.4
0.15x-0.07x = 2.4 - 1.4
0.08x = 1
x = 1/0.08 = 12.8 Gallons
Ans :
12.8 Gallons of 15% alcohol
7.2 Gallons of 7% alcohol